# Diffusion MRI with Machine Learning

### v1

@@ Abstract
Diffusion-weighted magnetic resonance imaging (dMRI) offers unique capabilities such as noninvasive assessment of brain’s micro-structure and structural connectivity. However, analyzing the dMRI data to extract useful information for clinical and scientific purposes is challenging. The dMRI measurements often suffer from strong noise and artifacts, there is usually high inter-session and inter-scanner heterogeneity in the data and considerable inter-subject variability in brain structure, and the relationship between measurements and the phenomena of interest can be highly complex. Recent years have witnessed increasing use of machine learning methods for dMRI analysis. This manuscript aims to assess these efforts, with a focus on methods that have addressed micro-structure mapping, tractography, white matter tract analysis, as well as data preprocessing and harmonization. We summarize the main findings, strengths, and weaknesses of the existing methods and suggest topics for future research. We find that machine learning may be exceptionally suited to tackle some of the difficult tasks in dMRI analysis. However, for this to happen, several shortcomings of existing methods and critical unresolved issues need to be addressed. These include deficient evaluation practices, lack of rich training datasets and validation benchmarks, as well as model generalizability, reliability, and explainability concerns.
@@ I Introduction
I-A Background, motivation, and the scope of this article
@@
Diffusion-weighted Magnetic Resonance Imaging (dMRI) is a widely used medical imaging modality. It has a unique role in neuroimaging, where it stands as the only noninvasive method for probing the micro-structural makeup and structural connectivity of the brain. It has facilitated the study of normal brain development and quantitative characterization of the impact of diseases and disorders such as epilepsy, multiple sclerosis (MS), and schizophrenia. As a result, dMRI is an indispensable tool in medicine and neuroscience, and it has been a major component of large neuroimaging initiatives.
@@
Raw dMRI data suffer from a host of imperfections and artifacts. Yet, these data need to be analyzed to uncover subtle differences or minute changes that reflect the underlying normal or abnormal variations in brain development. These neurodevelopmental processes, in turn, are highly complex, heterogeneous, and multi-factorial. Consequently, development and validation of computational methods for dMRI analysis are very challenging. Accurate and reproducible processing of dMRI data has been a long-standing challenge, and thousands of research papers have been devoted to addressing its various aspects. Several computational pipelines and software projects have aimed at standardizing and streamlining some of the more routine dMRI computations [1, 2, 3]. However, there are many persistent challenges that have not been fully addressed and there is an urgent need for new methods to enable higher accuracy, reproducibility, reliability, and computational speed for dMRI analysis [4, 5, 6].
@@
Classical dMRI analysis methods have been mostly based on conventional signal processing, biophysical modeling, and numerical optimization techniques. At the same time, some studies have advocated for machine learning and data-driven techniques. Over the past few years, these techniques have become more popular than before. This trend has been driven by more powerful machine learning models (mostly based on deep neural networks) and a greater appreciation of the power and flexibility of these methods. Overall, these methods have been shown to possess the potential to improve the speed, accuracy, and reproducibility of various computations in dMRI analysis. It appears that the recent works on this topic herald a new generation of dMRI analysis methods that may significantly complement, if not at least in some cases entirely replace, the more conventional techniques.
@@
Therefore, we think it is good time to review, summarize, and critically assess the main achievements of these works, highlight some of their major shortcomings and limitations, and point out future work that may contribute to this new field in dMRI. This manuscript focuses almost entirely on applications in neuroimaging. It deals with the processing steps after the reconstruction of the so-called q-space data. Although it briefly discusses data pre-processing steps such as denoising and artifact correction, it is mostly concerned with the more downstream computations such as white matter micro-structure estimation and tract-specific analysis.
I-B The challenging nature of dMRI analysis
@@
A primary source of difficulty in processing and analyzing the dMRI data is strong noise and artifacts [4, 7]. Measurement noise can significantly alter the analysis results in real-world applications [8, 9]. The echo planar imaging method that is used to acquire dMRI measurements can give rise to eddy-current- and susceptibility-induced distortions. Subject motion during data acquisition is another persistent challenge in MRI. These factors can significantly impact the accuracy of quantitative micro-structure mapping and structural connectivity analysis [10, 11, 12]. Furthermore, the relationship between the dMRI measurements and the underlying micro-structure of neuronal tissue is very complex [13, 14]. Another source of difficulty is the measurement requirements of advanced dMRI models. Estimation of multi-compartment models of tissue micro-structure and of complex white matter fiber configurations requires densely-sampled q-space measurements. Obtaining such measurements would require long scan times that may not be affordable in many settings. Even with reasonably large number of measurements, standard computational methods can produce erroneous results [15]. As an example, conventional linear and non-linear least-squares techniques for diffusion kurtosis imaging tend to produce physically implausible results [16, 17]. Higher-level computations such as streamline tractography and tract-specific analysis suffer from various difficulties that have been the subject of intense research for the past two decades [18, 5, 19, 20].
I-C Why machine learning may help
@@
Machine learning methods encompass a rich set of flexible and powerful approaches that have the potential to improve upon the more conventional techniques for dMRI analysis. The advantages of machine learning methods over classical techniques depend on the specific tasks, which are discussed in more detail in the following sections. Here, we list some of the main reasons why machine learning may be useful for dMRI analysis.
@@
  • Machine learning models can represent highly complex functions. For neural networks, universal approximation theorems state that they can represent all functions of practical interest [23]. This means that machine learning models are far less restricted than conventional methods in terms of the complexity of the phenomena that they can represent. This can be a significant advantage because the true mapping between the input and output may be outside the scope of the mathematical models imposed by conventional analysis methods [24, 25].
  • As a result of the above properties, machine learning methods can simplify the complex multi-stage pipelines into simpler end-to-end methods. Many dMRI analysis tasks such as estimation of tissue micro-structure or segmentation of white matter tracts rely on a sequence of computations that are optimized separately and independently [26, 27]. With machine learning methods, it is often possible to combine these into a single model that can be optimized jointly [27].
  • Many machine learning models can effectively and seamlessly integrate various inputs and sources of prior knowledge such as spatial information or other MRI contrasts. As an example, recent works have shown that deep learning methods can easily leverage anatomical MRI data on top of dMRI measurements to improve various dMRI analysis tasks including micro-structure estimation [28], super-resolution [29], tissue segmentation [26], and distortion correction [30]. Given the rich spatial regularity in many neuroimaging problems, incorporation of spatial information may significantly improve the model performance. While such prior knowledge may be too complex to mathematically formulate with conventional methods, machine learning models can effortlessly learn them directly from data. Modern neural networks such as Convolutional Neural Networks (CNNs), graph CNNs, and transformers have been shown to be especially well-suited for this purpose [31, 32, 28, 33].
@@
  • Machine learning models may be trained on data with various types and different levels of imperfections such as noise, motion, sub-optimal acquisition protocols, and imaging artifacts. This way, they may gain a degree of robustness with respect to these data imperfections by learning to factor them out in their computations [38, 36, 39, 40, 41, 32]. It has been reported, for example by Wegmayr et al. for tractography [42], that machine learning models trained on data with noisy labels may produce more accurate and less noisy predictions on test data. Weine et al. developed a pipeline for generating synthetic cardiac DTI data that incorporated simulated heart motion and showed that a deep learning model trained on such data outperformed standard DTI estimation methods [32].
@@
  • Finally, as a powerful data analysis method, machine learning may offer new ways of analyzing the data and provide new insights into the potential and limitations of dMRI that are difficult or impossible to achieve with conventional signal processing methods. A good case in point is the recent work of Cai et al. on tractography [44], where the authors have investigated the possibility of performing tractography based purely on anatomical MRI data without any diffusion encoding. Their results suggest that tractography and structural connectivity analysis based on anatomical MRI may be comparable with those based on dense dMRI measurements, raising questions about whether standard streamline tractography is driven by tissue micro-structure. Another example is the use of autoencoders or similar bottleneck networks to characterize the information content of dMRI signal and discover common features between different biophysical models [45, 46].

### v2

@@ Abstract
Diffusion-weighted magnetic resonance imaging (dMRI) offers unique capabilities including noninvasive probing of brain’s tissue microstructure and structural connectivity. It is widely used for clinical assessment of brain pathologies and for neuroscience research. Analyzing the dMRI data to extract useful information for medical and scientific purposes can be challenging. The dMRI measurements often suffer from strong noise and artifacts, there is usually high inter-session and inter-scanner variability in the data, and considerable inter-subject heterogeneity in brain structure. Moreover, the relationship between measurements and the phenomena of interest can be highly complex. Recent years have witnessed increasing use of machine learning methods for dMRI analysis. This manuscript aims to assess these efforts, with a focus on methods that have addressed data preprocessing and harmonization, microstructure mapping, tractography, and white matter tract analysis. We study the main findings, strengths, and weaknesses of the existing methods and suggest topics for future research. We find that machine learning may be exceptionally suited to tackle some of the difficult tasks in dMRI analysis. However, for this to happen, several shortcomings of existing methods and critical unresolved issues need to be addressed. These include deficient evaluation practices, lack of rich training datasets and validation benchmarks, as well as model generalizability, reliability, and explainability concerns.
Index Terms:
Diffusion MRI, Machine Learning, Artificial Intelligence, Deep Learning
@@ I Introduction
I-A Background and motivation
@@
Diffusion-weighted Magnetic Resonance Imaging (dMRI) is a widely used medical imaging modality [1, 2]. It has a unique role in neuroimaging, where it stands as the only noninvasive method for probing the tissue microstructural makeup and structural connectivity of the brain [3, 4]. It has facilitated the study of normal brain development and quantitative characterization of the impact of diseases and disorders [5, 6]. As a result, dMRI has become an indispensable tool in medicine and neuroscience, and it has been a major component of large neuroimaging initiatives [7, 8].
@@
Raw dMRI data can suffer from a host of imperfections and artifacts [9, 10]. Yet, these data need to be analyzed to uncover subtle differences or minute changes that reflect the underlying normal or abnormal variations in brain development. These neurodevelopmental processes, in turn, are highly complex, heterogeneous, and multi-factorial. Consequently, development and validation of computational methods for dMRI analysis are difficult. Accurate and reproducible processing of dMRI data has been a long-standing challenge, and thousands of research papers have been devoted to addressing its various aspects. Several computational pipelines and software projects have aimed at standardizing and streamlining some of the more routine dMRI computations [11, 12, 13]. However, there are many persistent challenges that have not been fully addressed and there is an urgent need for new methods to enable higher accuracy, reproducibility, reliability, and computational speed [9, 14, 15].
@@
Classical dMRI analysis methods have been mostly based on conventional signal processing, biophysical modeling, and numerical optimization techniques. Meanwhile, many studies have advocated for machine learning and data-driven techniques [16, 17, 18]. These techniques have become more popular in recent years. This trend has been driven by more powerful machine learning models (mostly based on deep neural networks) and a greater appreciation of the power and flexibility of these methods. Overall, these methods have been shown to possess the potential to improve the speed, accuracy, and reproducibility of various computations in dMRI analysis such as data pre-processing [19] and harmonization [20], tissue microstructure mapping [16, 21, 22], tractography [18], tract-specific analysis [17], and population studies [23]. It appears that recent studies herald a new generation of dMRI analysis methods that may significantly complement, if not at least in some cases entirely replace, the more conventional techniques. Therefore, it is good time to review, summarize, and critically assess the achievements of these works, highlight their shortcomings and limitations, and point out future work that may contribute to this new field in dMRI.
I-B Scope and organization of this manuscript
@@
This manuscript focuses almost entirely on applications in neuroimaging. It deals with the processing steps after the reconstruction of the so-called q-space data. It is concerned with data pre-processing steps such as denoising and artifact correction as well as downstream computations such as white matter microstructure mapping and tract-specific analysis.
The search to find the relevant works to be included in this paper was conducted in PubMed and Google Scholar. The initial search was performed in June 2023 and then repeated in June 2024. The search terms included “diffusion MRI” and “machine learning” or “deep learning”. More detailed searches were performed in reputable journals (e.g., Medical Image Analysis and IEEE Transactions on Medical Imaging) and conferences (e.g., MICCAI). Given the wide scope of this paper, we excluded the works that had little methodological novelty or limited experimental validation. Additionally, where two or more papers presented similar ideas and methods, we chose the paper with more comprehensive experimental results.
This monograph starts with a brief description of challenges in dMRI data analysis (Section II) and general overview of the reasons why machine learning may be effective in addressing these challenges (Section III). Section IV reviews prior works that have used machine learning in dMRI. The main applications considered include data preprocessing and quality enhancement, data harmonization, estimation of tissue microstructure and fiber orientation, tractography, tract-specific analysis, registration and segmentation. Figure 1 shows the outline of that section. Following that, Section V discusses the main technical considerations, challenges, and open questions. Some of the main topics discussed in that section include validation approaches, inherent limitations of machine learning, data, ground truth, and modeling considerations, as well as model explainability, uncertainty, and reliability concerns. Figure 2 shows the outline of that section. A short Conclusion section will present the closing remarks.
II The challenging nature of dMRI analysis
@@
The challenges and pitfalls of analyzing the dMRI data have been discussed in dedicated publications [9, 15, 24]. A primary source of difficulty in dMRI analysis is measurement noise and artifacts [9, 25]. Measurement noise can significantly alter the analysis results even in relatively simple computations such as diffusion tensor estimation [26, 27]. It can be difficult to suppress the noise and artifacts due to their complex distributions. The echo planar imaging method that is used to acquire dMRI measurements can give rise to eddy-current- and susceptibility-induced distortions. Subject motion during data acquisition is another persistent challenge in MRI. These factors can significantly impact the accuracy of quantitative tissue microstructure mapping and structural connectivity analysis [28, 29, 30]. Furthermore, the relationship between the dMRI measurements and the underlying microstructure of neuronal tissue is complex [31, 32]. Another source of difficulty is the measurement requirements of advanced dMRI models. Estimation of multi-compartment models of tissue microstructure and of complex white matter fiber configurations requires densely-sampled q-space data. Obtaining such measurements would require long scan times that may not be affordable in many settings. Even with a reasonably large number of measurements, standard computational methods can produce erroneous results [33]. As only one example, conventional linear and non-linear least-squares techniques for diffusion kurtosis imaging (DKI [34]) tend to produce physically implausible results [35, 36]. Streamline tractography and tract-specific analysis face inherent challenges and ambiguities such as fiber crossings and bottlenecks [24, 37, 38, 39], which can give rise to high false positive rates and erroneous results. Cross-subject comparisons and population studies can also be significantly hampered by high inter-scanner data heterogeneity [40, 41]. Furthermore, most applications are also challenged by paucity of ground truth and lack of universally accepted performance metrics.
III Why machine learning may help
@@
Machine learning encompasses a rich set of flexible and powerful methods that have the potential to improve upon the more conventional techniques for dMRI analysis. The advantages of machine learning methods over classical techniques depend on the specific tasks, which are discussed in more detail in the following sections. Here, we list some of the main reasons why machine learning may be useful for dMRI analysis.
@@
  • Machine learning models can represent highly complex functions. For neural networks, universal approximation theorems state that they can represent all functions of practical interest [43]. This means that machine learning models are far less restricted than conventional methods in terms of the complexity of the phenomena that they can represent. This can be a significant advantage because the true mapping between the input and output may be outside the scope of the mathematical models imposed by conventional methods [44, 45].
  • As a result of the above properties, machine learning methods can simplify the complex multi-stage pipelines into simpler end-to-end methods. Many dMRI analysis tasks such as estimation of tissue microstructure or segmentation of white matter tracts rely on a sequence of computations that are optimized separately and independently [16, 17]. With machine learning methods, it is often possible to combine these into a single model that can be optimized jointly [17].
  • Many machine learning models can effectively and seamlessly integrate various inputs, constraints, and sources of prior knowledge, such as other MRI contrasts or spatial information. As an example, recent works have shown that deep learning methods can easily leverage anatomical MRI data (on top of dMRI data) to improve microstructure estimation [46], super-resolution [47], tissue segmentation [16], and distortion correction [19]. Given the rich spatial regularity in many neuroimaging problems, incorporation of spatial information may significantly improve the model performance. Spatial information can be too complex to mathematically formulate with conventional methods and, if not done properly, can have a negative impact on performance [48]. Machine learning models, on the other hand, can effortlessly learn this information directly from data. Modern neural networks such as Convolutional Neural Networks (CNNs), graph CNNs, and transformers have been shown to be especially well-suited for this purpose [49, 50, 46, 51].
@@
  • Machine learning models may be trained on data with various types and different levels of imperfections such as noise, motion, sub-optimal acquisition protocols, and imaging artifacts. This way, they may gain a degree of robustness with respect to these data imperfections by learning to factor them out in their computations [55, 21, 56, 57, 58, 50]. This can be a unique advantage since mathematically modeling these data imperfections can be difficult or impossible. It has been reported, for example by Wegmayr et al. for tractography [59], that machine learning models trained on data with noisy labels may produce more accurate and less noisy predictions on test data. Weine et al. developed a pipeline for generating synthetic cardiac diffusion tensor imaging (DTI) data that incorporated simulated heart motion and showed that a deep learning model trained on such data outperformed standard DTI estimation methods [50].
@@
  • By offering new ways of analyzing the dMRI data, machine learning methods provide new insights into the potential and limitations of this imaging modality. A good case in point is the recent work of Cai et al. on tractography [61], where the authors have investigated the possibility of performing tractography based purely on anatomical MRI data without any diffusion encoding. Their results suggest that tractography and structural connectivity analysis based on anatomical MRI may be comparable with those based on dMRI measurements, raising questions about whether standard streamline tractography is driven by tissue microstructure. Other works have used autoencoders and similar bottleneck networks to characterize the information content of dMRI signal and discover common features between different biophysical models [62, 63].
IV Prior works on dMRI analysis with machine learning